OpenStudy (anonymous):

HOW TO SOLVE THIS STUPID THING... Solve for y: x=(y-7/2)^1/5+3

OpenStudy (gorv):

x-3=(y-7/2)^1/5 raise both side by power 5 (x-3)^5=y-7/2 (x-3)^5+7/2=y

OpenStudy (anonymous):

Is it ^(1/5+3) or ^(1/5) + 3

OpenStudy (anonymous):

How did my question get so many viewers???? OH MY GAWD HALP!

OpenStudy (anonymous):

You see these questions have an actual value rather than the previous spam you post.

OpenStudy (usukidoll):

\[x =( y-\frac{ 7 }{ 2 })^{\frac{ 1 }{ 5 }}+3\]

OpenStudy (anonymous):

:)

OpenStudy (gorv):

bcozz no new Q are there to solve

OpenStudy (anonymous):

\( \newcommand\permute[2]{^{#1}P_{#2}} \) Hmm

OpenStudy (anonymous):

mmh

OpenStudy (anonymous):

Wait a moment...

OpenStudy (gorv):

waitin.............................................

OpenStudy (anonymous):